OpenStudy (anonymous):

please help me everyone . what is the derivative of: f(t)=20[1-cos^2(t/1000pi)]

OpenStudy (abb0t):

20 is a constant and can be factored out. derivative of 1 is zero for derivative of: \[\cos^2(\frac{ t }{ 1000 \pi })\] use chain rule. first use the property: \[\frac{ d }{ dx }a^n = na^{n-1}\] where a is 'a' constant and 'n' is some numerical value. 'a' can also be some function treated as a constant. For instance, \[\frac{ d }{ dx }\sin^2(x) = 2\sin^{2-1}(x) \cos(x)\] the cos(x) is just the derivative of the inside function (chain rule).